# Uncertainty Quantification for LLMs through Minimum Bayes Risk: Bridging Confidence and Consistency

### v4

@@ D.1 Information-Based Methods
Maximum Sequence Probability (MSP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. In its simplest form it is given by (6). An equivalent (in terms of ordering of predictions for different inputs) but more numerically stable formulation is logarithmic. It is defined as:
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models [Fomicheva et al., 2020]. It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence [Fomicheva et al., 2020]. Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
@@
The TokenSAR method, introduced in Duan et al. [2024], generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function and token relevance function , the uncertainty estimate is calculated as:
@@ D.2 Consistency-Based Methods
The work Lin et al. [2024] offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: .
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model He et al. [2021], following Kuhn et al. [2023]. For each pair of sequences, an NLI model predicts two probabilities: , indicating entailment, and , indicating contradiction. Similarity is then defined as either or .
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian [Lin et al., 2024]. Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, . Two sentences are grouped into the same cluster if the following conditions are satisfied: . This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by .
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. The work Lin et al. [2024] proposes computing an averaged similarity matrix as . The Laplacian for the matrix is defined as , where is a (diagonal) degree matrix with elements . Consequently, the following formula is derived:
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, the work Lin et al. [2024] proposes to use the diagonal Degree Matrix which represents the total similarity of each answer with all others. The corrected trace of provides an average pairwise distance between answers, and uncertainty is computed as:
@@ D.3 Information-Based Methods with Repeated Sampling
We can compute the entropy on the sequence level , where the expectation is taken over the sequences randomly generated from the distribution . Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences via random sampling and compute Monte Carlo Sequence Entropy [Kuhn et al., 2023]:
@@
Semantically weighted averaging. Semantic Entropy Kuhn et al. [2023] addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (25). The method clusters generated sequences into semantically homogeneous groups (where ) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
@@
SentenceSAR Duan et al. [2024] enhances the probability of sentences that are more relevant. It uses a sentence relevance measure to evaluate the relevance of with respect to . SentenceSAR is calculated as:
@@
The combination of SentenceSAR and TokenSAR results in a unified method called SAR Duan et al. [2024]. In this approach, the generative probability in the SentenceSAR formula is replaced with the token-shifted probability , creating a comprehensive measure that integrates both sentence- and token-level adjustments.
The aggregation approaches like Semantic Entropy Kuhn et al. [2023] or SAR Duan et al. [2024] can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
@@ Appendix F Alternative Performance Metrics for PRR
Tables 14 and 15 report PRRs for these metrics on all models for NMT and QA tasks. MetricX Juraska et al. [2024] was used for NMT and GPT-as-a-judge (gpt-4o-2024-08-06) was used to score QA datasets. The following is the prompt used to facilitate GPT QA scoring:

### v5

@@ C.1 Choice of Similarity Function
AlignScore [zha2023alignscore] with AlignScore-large model;
RougeL [lin-2004-rouge];
NLI [he2021deberta] based on microsoft/deberta-large-mnli model;
CrossEncoder [liu2019robertarobustlyoptimizedbert] based on cross-encoder/stsb-roberta-large model.
@@ D.1 Information-Based Methods
Sequence Probability (SP) is one of the simplest and most direct methods for estimating uncertainty. It measures the probability of the most likely output sequence given a specific input. Thus, uncertainty is quantified by calculating the probability of the sequence with the highest likelihood, under the assumption that the model is most confident in this output. In its simplest form it is given by (6). An equivalent (in terms of ordering of predictions for different inputs) but more numerically stable formulation is logarithmic. It is defined as:
Perplexity (PPL) is another widely used metric for estimating uncertainty in language models [fomicheva-etal-2020-unsupervised]. It measures the model’s confidence by evaluating the average likelihood of generating the sequence tokens:
Mean Token Entropy takes a broader view of uncertainty by considering the token-level predictions across the entire sequence [fomicheva-etal-2020-unsupervised]. Instead of evaluating the model’s confidence in a single output or individual token predictions, Mean Token Entropy calculates the average entropy of the token probability distributions for each token in the sequence:
@@
The TokenSAR method, introduced in [duan-etal-2024-shifting], generalizes length-normalized log probability by computing a weighted average of the negative log probabilities of generated tokens, where weights are based on token relevance to the overall text. Using a similarity function and token relevance function , the uncertainty estimate is calculated as:
@@ D.2 Consistency-Based Methods
The work [lin2023generating] offers two similarity measures to evaluate the similarity of sequences. The first is the Jaccard similarity, which treats sequences as sets of words and calculates the proportion of shared words to the total number of unique words in both sequences: .
Natural Language Inference (NLI) provides another method for computing similarity between sequences. We use the DeBERTa-large NLI model [he2021deberta], following [kuhn2023semantic]. For each pair of sequences, an NLI model predicts two probabilities: , indicating entailment, and , indicating contradiction. Similarity is then defined as either or .
Among the simplest consistency-based approaches are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian [lin2023generating]. Number of Semantic Sets estimates how many distinct “meanings” the model produces by clustering its outputs with an NLI model. The number of semantic sets is initially equal to the total number of generated answers, . Two sentences are grouped into the same cluster if the following conditions are satisfied: . This computation is performed for all pairs of answers, and the final number of distinct clusters is denoted by .
Sum of Eigenvalues of the Graph Laplacian examines global diversity: it constructs a similarity matrix among the sampled outputs and computes a continuous uncertainty score from the eigenvalues of the Laplacian of that similarity graph. The work [lin2023generating] proposes computing an averaged similarity matrix as . The Laplacian for the matrix is defined as , where is a (diagonal) degree matrix with elements . Consequently, the following formula is derived:
Both Number of Semantic Sets and Sum of Eigenvalues of the Graph Laplacian effectively capture overall variation in generated text but cannot produce an individual uncertainty score for each output. To address this, the work [lin2023generating] proposes to use the diagonal Degree Matrix which represents the total similarity of each answer with all others. The corrected trace of provides an average pairwise distance between answers, and uncertainty is computed as:
@@ D.3 Information-Based Methods with Repeated Sampling
We can compute the entropy on the sequence level , where the expectation is taken over the sequences randomly generated from the distribution . Unfortunately, while for token level, we have an exact way of computing the entropy, for the sequence level, we need to adhere to some approximations. In practice, we can use Monte-Carlo integration, i.e. generate several sequences via random sampling and compute Monte Carlo Sequence Entropy [kuhn2023semantic]:
@@
Semantically weighted averaging. Semantic Entropy [kuhn2023semantic] addresses the issue of generated sequences with similar meanings but differing probabilities according to the model, which can heavily influence the resulting entropy value (25). The method clusters generated sequences into semantically homogeneous groups (where ) using a bi-directional entailment algorithm. Probabilities of sequences are averaged within each cluster. The entropy estimate is then defined as:
@@
SentenceSAR [duan-etal-2024-shifting] enhances the probability of sentences that are more relevant. It uses a sentence relevance measure to evaluate the relevance of with respect to . SentenceSAR is calculated as:
@@
The combination of SentenceSAR and TokenSAR results in a unified method called SAR [duan-etal-2024-shifting]. In this approach, the generative probability in the SentenceSAR formula is replaced with the token-shifted probability , creating a comprehensive measure that integrates both sentence- and token-level adjustments.
The aggregation approaches like Semantic Entropy [kuhn2023semantic] or SAR [duan-etal-2024-shifting] can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
@@
D.4 Verbalized Uncertainty
Verbalized uncertainty methods refer to approaches that prompt a model to explicitly express its confidence. In our experiment, we utilize the P(True) method, implemented following the description by [kadavath2022language]. Specifically, the model is presented with the original question and its answer, and then prompted to indicate whether the answer is True or False. We use the negative log-probability of the token “True” as the uncertainty score.
@@ Appendix F Alternative Performance Metrics for PRR
Tables 14 and 15 report PRRs for these metrics on all models for NMT and QA tasks. MetricX [juraska-etal-2024-metricx] was used for NMT and GPT-as-a-judge (gpt-4o-2024-08-06) was used to score QA datasets. The following is the prompt used to facilitate GPT QA scoring:
@@ Appendix H Computational Budget
Appendix I CoCoA Light Training Details
The model architecture consisted of a simple multilayer perceptron (MLP) with a single hidden layer:
Input dimension: equal to the embedding size of the corresponding base model (4096 for Mistral7b-Base and Llama8b-Base, 3072 for Falcon7b-Base).
Hidden dimension: 2048.
Output dimension: 4096.
Dropout: 0.1 applied between layers.
We trained the MLP on mean pooled hidden-layer embeddings extracted from the middle layer of an LLM. Table 18 details the size of a train set for each model and dataset.
| Dataset | LLaMA | Falcon | Mistral |
| CoQA | 10,000 | 10,000 | 10,000 |
| GSM8K | 3,000 | 3,000 | 2,500 |
| MMLU | 1,461 | 1,461 | 1,461 |
| TriviaQA | 10,000 | 10,000 | 10,000 |
| WMT14 Fr-En | 6,000 | 6,000 | 6,000 |
| WMT19 De-En | 6,000 | 6,000 | 6,000 |
| XSum | 7,500 | 5,000 | 6,500 |
Training was conducted for 20 epochs with the following hyperparameters:
Batch size: 4 (per device) with gradient accumulation of 7 steps (effective batch size of 28).
Learning rate: with AdamW optimizer.
Weight decay: 0.1.
Warmup ratio: 0.05.
Gradient clipping: 1.0.
Appendix J Larger Model Experiments
We additionally evaluate our method using the Gemma 3 12B-Base model to assess its performance on a larger-scale architecture. As shown in Table 19, CoCoA continues to achieve the best average performance among all methods.
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| MCSE | -0.014 | 0.277 | 0.408 | 0.262 | 0.556 | 0.449 | 0.386 |
| MCNSE | -0.031 | 0.393 | 0.442 | 0.280 | 0.595 | 0.468 | 0.344 |
| Semantic Entropy | -0.011 | 0.273 | 0.430 | 0.287 | 0.621 | 0.531 | 0.407 |
| DegMat | 0.130 | 0.246 | 0.370 | 0.356 | 0.692 | 0.561 | 0.359 |
| EigValLaplacian | 0.130 | 0.195 | 0.307 | 0.323 | 0.667 | 0.526 | 0.312 |
| SAR | 0.087 | 0.429 | 0.494 | 0.335 | 0.688 | 0.582 | 0.418 |
| P(True) | -0.022 | 0.012 | -0.020 | -0.006 | 0.195 | 0.046 | 0.145 |
| Consistency Light | -0.006 | 0.328 | 0.483 | 0.345 | 0.709 | 0.408 | 0.421 |
| Consistency | -0.163 | 0.452 | 0.496 | 0.311 | 0.657 | 0.293 | 0.387 |
| MSP | 0.257 | 0.315 | 0.499 | 0.319 | 0.672 | 0.630 | 0.335 |
| 0.308 | 0.424 | 0.641 | 0.372 | 0.723 | 0.617 | 0.391 | |
| Light | 0.287 | 0.416 | 0.634 | 0.356 | 0.720 | 0.613 | 0.360 |
| PPL | 0.303 | 0.369 | 0.451 | 0.288 | 0.681 | 0.630 | 0.275 |
| 0.339 | 0.427 | 0.539 | 0.362 | 0.728 | 0.617 | 0.439 | |
| Light | 0.327 | 0.489 | 0.518 | 0.334 | 0.727 | 0.613 | 0.353 |
| MTE | 0.284 | 0.372 | 0.413 | 0.269 | 0.664 | 0.617 | 0.310 |
| 0.326 | 0.422 | 0.527 | 0.344 | 0.728 | 0.592 | 0.468 | |
| Light | 0.308 | 0.501 | 0.498 | 0.313 | 0.719 | 0.565 | 0.386 |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| MCSE | 0.051 | 0.346 | 0.447 | 0.273 | 0.531 | 0.447 | 0.495 |
| MCNSE | 0.088 | 0.441 | 0.456 | 0.319 | 0.580 | 0.465 | 0.533 |
| Semantic Entropy | 0.054 | 0.348 | 0.492 | 0.294 | 0.597 | 0.527 | 0.505 |
| DegMat | 0.165 | 0.294 | 0.372 | 0.366 | 0.659 | 0.558 | 0.457 |
| EigValLaplacian | 0.164 | 0.228 | 0.306 | 0.327 | 0.634 | 0.522 | 0.415 |
| SAR | 0.109 | 0.474 | 0.526 | 0.353 | 0.661 | 0.577 | 0.593 |
| P(True) | -0.029 | 0.037 | -0.015 | 0.019 | 0.193 | 0.046 | 0.278 |
| Consistency Light | 0.076 | 0.478 | 0.631 | 0.373 | 0.671 | 0.443 | 0.744 |
| MSP | 0.188 | 0.353 | 0.466 | 0.361 | 0.646 | 0.624 | 0.223 |
| 0.251 | 0.542 | 0.718 | 0.406 | 0.693 | 0.613 | 0.433 | |
| PPL | 0.178 | 0.501 | 0.619 | 0.349 | 0.660 | 0.624 | 0.750 |
| 0.211 | 0.576 | 0.692 | 0.400 | 0.700 | 0.613 | 0.805 | |
| MTE | 0.168 | 0.431 | 0.487 | 0.295 | 0.643 | 0.612 | 0.703 |
| 0.212 | 0.546 | 0.641 | 0.369 | 0.702 | 0.591 | 0.796 | |
| Method | Dataset | ||||||
| XSUM | WMT14FrEn | WMT19DeEn | CoQA | TriviaQA | MMLU | GSM8k | |
| MCSE | 0.120 | 0.290 | 0.361 | 0.271 | 0.544 | 0.487 | 0.493 |
| MCNSE | 0.126 | 0.366 | 0.439 | 0.308 | 0.597 | 0.497 | 0.464 |
| Semantic Entropy | 0.121 | 0.267 | 0.396 | 0.307 | 0.611 | 0.550 | 0.468 |
| DegMat | 0.195 | 0.222 | 0.318 | 0.399 | 0.684 | 0.571 | 0.373 |
| EigValLaplacian | 0.197 | 0.170 | 0.254 | 0.367 | 0.653 | 0.522 | 0.319 |
| SAR | 0.204 | 0.409 | 0.464 | 0.372 | 0.680 | 0.591 | 0.508 |
| P(True) | 0.030 | 0.043 | 0.018 | 0.014 | 0.204 | 0.079 | 0.313 |
| Consistency Light | 0.192 | 0.311 | 0.419 | 0.368 | 0.690 | 0.421 | 0.467 |
| MSP | 0.222 | 0.390 | 0.491 | 0.240 | 0.330 | 0.542 | 0.441 |
| 0.282 | 0.452 | 0.598 | 0.347 | 0.624 | 0.574 | 0.498 | |
| PPL | 0.187 | 0.400 | 0.526 | 0.183 | 0.344 | 0.542 | 0.468 |
| 0.270 | 0.476 | 0.587 | 0.337 | 0.638 | 0.574 | 0.557 | |
| MTE | 0.139 | 0.425 | 0.515 | 0.251 | 0.628 | 0.623 | 0.498 |
| 0.230 | 0.451 | 0.528 | 0.361 | 0.712 | 0.610 | 0.575 | |